# Repeatable and Reliable Efforts of Accelerated Risk Assessment in Robot Testing

### v1

@@ 2.2 Sampling-based risk estimate
6:    $r=\Big{(}\sum_{j=1}^{i}(1-\delta_{T_{i}\cap\mathcal{C},\emptyset})\cdot p_{s}(x)\Big{)}/i$
@@
6:    $r=\Big{(}\sum_{j=1}^{i}(1-\delta_{T_{i}\cap\mathcal{C},\emptyset})\cdot p(x_{i})/q(x_{i})\Big{)}/i$
@@
Moreover, the variance-based criterion and RHW form only a necessary condition of good performance. The threshold $s_{r}$ is typically fixed for all testing attempts and testing subjects. When RHW is used as the basis of a runtime termination observer, one could end up with multiple different risk estimates, all satisfying the given threshold (i.e. the repeatability problem). More importantly, these different estimates could come from substantially different testing efforts. The number of tests required could also fluctuate widely as one tests different functions $f\in S^{S\times U\times W}$. This variability introduces further complications, as there is no clear way to anticipate the extent of testing efforts or their variability across different test subjects. As a result, the previously mentioned two algorithms fail to meet the criteria to be $\gamma$-reliable per Definition 2 to be presented in the next section.
@@ 2.3 Repeatable & reliable risk estimate
[$\beta$-Repeatability] A sampling-based risk-estimate testing algorithm $\mathcal{TE}$ as defined in Algorithm 1 and Algorithm 2 is $\beta$-repeatable if for some $\beta\in(0,1)$ and any two executions of the same algorithm with all other configurations following the same level of randomness, $\mathcal{TE}^{i}$, $\mathcal{TE}^{j}$ and $i\neq j$,
@@
[$\gamma$-Reliability] Let $T_{jk}$ denote the set of state trajectories collected from the $j$-th execution of a testing algorithm, $\mathcal{TE}^{j}$, with a certain testing system $f_{k}\in S^{S\times U\times W}$. A sampling-based risk-estimate testing algorithm $\mathcal{TE}$ as defined in Algorithm 1 and Algorithm 2 is $\gamma$-reliable if, with all other configurations following the same level of randomness, it is $\beta$-repeatable for some $\beta\in(0,1)$, and there exists $\gamma\in\mathbb{Z}$, $\mathcal{F}\subset S^{S\times U\times W}$, and
@@ 3.1 The main algorithm, theorem, and proof
Let $\phi_{f}:X\rightarrow\mathbb{B}$ be the failure check function (which composes the testing system dynamics (1), state trajectory collection, and the intersection check against $\mathcal{C}$) with a boolean output (w.l.o.g., let $\phi_{f}(x)\in\{0,1\},\forall x$). We have $r^{*}=\mathbb{E}_{p}\phi(x)$ and the risk estimate through Algorithm 2 as $r=\frac{1}{n}\sum_{i=1}^{n}\phi_{f}(x_{i})p(x_{i})/q(x_{i})$. Let $L=D(p\mid\mid q)$ We have for $n=e^{L+t}$ with some $t\in\mathbb{R}_{\geq 0}$,
@@
$\leq\sqrt{\mathbb{E}(\phi_{f}(x)^{2})}\cdot\Bigg{(}e^{-\frac{c}{4}}+2\sqrt{\mathbb{P}\Big{(}\log\frac{p(x)}{q(x)}>L+\frac{c}{2}\Big{)}}\Bigg{)}$
$\leq\bar{r}\cdot\Bigg{(}e^{-\frac{c}{4}}+2\sqrt{\mathbb{P}\Big{(}\log\frac{p(x)}{q(x)}>L+\frac{c}{2}\Big{)}}\Bigg{)}\leq\tau^{\prime}.$
@@
With $|r-r^{*}|\leq 1$, and by Hoeffding’s inequality, we further have $\mathbb{P}(|r-r^{*}|\geq\tau)\leq 2e^{-2n\tau^{2}}$. This proves the $\beta$-repeatability of estimation accuracy $\tau$ except with probability $2e^{-2n\tau^{2}}$. Furthermore, for $\gamma$-reliability to be satisfied, the number of samples $n$ needs to be upper bounded by $\gamma$. It is thus expected that the importance distribution $q$ satisfying:
@@ 3.2 Discussions
Moreover, the nominal and importance distributions are $p$ and $q$, respectively. Note in the general statistics applications, $p$ and $q$ may not be known properties [3], hence the presented results may not be as useful or applicable. However, in the context of testing, $p$ is often expected as part of the specifications made by standard organizations and regulatory authorities. $q$ is required for testing execution by operators and testing agencies.
@@
Not every importance sampling based testing algorithm is an accelerated testing algorithm. The accelerated performance largely relies on the properly chosen importance distribution $q$, which is out of the scope of this paper [29, 28, 10, 13, 12]. The proposed results in this section are applicable for an arbitrary pair of target and importance distributions $(p,q)$.
@@
The reliability of Algorithm 3, revealed by Theorem 3.1, is jointly determined by Kullback–Leibler (KL) divergence between the two distributions and the value of $c$ which is mainly associated with the typical order of fluctuations of $\log(p(X)/q(X))$ from its expected value. In practice, the KL-divergence between $p$ and $q$ is immediate given the above, and one can numerically approximate the minimum estimate of $c$ that satisfies (4) (e.g., an incremental trial-and-error search of $c$ values from zero as demonstrated in line 4-7 in Algorithm 3 or some other bisection alternatives). Note $c_{r}$ is preferably a small value to enhance the efficiency, but the search with large steps will not undermine the theoretical properties dictated from Theorem 3.1. These two values (order of fluctuations and KL-divergence) critically influence the necessary testing efforts to minimize the risk estimation error (subject to the $\ell_{1}$-norm bound of $\tau$) with a high probability of $1-2e^{-2n\tau^{2}}$, thereby insuring $\gamma$-reliability (and $\beta$-repeatability).

### v2

@@ 1.1 Accelerated testing for risk estimate
Remark 1
This study focuses on importance sampling based testing algorithms. While other testing algorithms for falsification [21, 1], verification [10, 20], and validation [19, 30] may share similar concerns of repeatability and reliability, they are out of the scope of this paper.
@@ 2.2 Sampling-based risk estimate
6:    $r=\Big{(}\sum_{j=1}^{i}(1-\delta_{T_{i}\cap\mathcal{C},\emptyset})\cdot p_{s}(x_{j})\Big{)}/i$
@@
6:    $r=\Big{(}\sum_{j=1}^{i}(1-\delta_{T_{j}\cap\mathcal{C},\emptyset})\cdot p(x_{j})/q(x_{j})\Big{)}/i$
@@
Moreover, the variance-based criterion and RHW form only a necessary condition for good performance. The threshold $s_{r}$ is typically fixed for all testing attempts and testing subjects. When RHW is used as the basis of a runtime termination observer, one could end up with multiple different risk estimates, all satisfying the given threshold (i.e. the repeatability problem). More importantly, these different estimates could come from substantially different testing efforts. The number of tests required could also fluctuate widely as one tests different functions $f\in S^{S\times U\times W}$. This variability introduces further complications, as there is no clear way to anticipate the extent of testing efforts or their variability across different test subjects. As a result, the previously mentioned two algorithms fail to meet the criteria to be $\gamma$-reliable per Definition 2 to be presented in the next section.
@@ 2.3 Repeatable & reliable risk estimate
[$\beta$-Repeatability] A sampling-based risk-estimate testing algorithm $\mathcal{TE}$ (e.g. Algorithm 1 and Algorithm 2) is $\beta$-repeatable if for some $\beta\in(0,1)$ and any two executions of the same algorithm with all other configurations following the same level of randomness, $\mathcal{TE}^{i}$, $\mathcal{TE}^{j}$ and $i\neq j$,
@@
[$\gamma$-Reliability] Let $T_{jk}$ denote the set of state trajectories collected from the $j$-th execution of a testing algorithm, $\mathcal{TE}^{j}$, with a certain testing system $f_{k}\in S^{S\times U\times W}$. A sampling-based risk-estimate testing algorithm $\mathcal{TE}$ (e.g. Algorithm 1 and Algorithm 2) is $\gamma$-reliable if, with all other configurations following the same level of randomness, it is $\beta$-repeatable for some $\beta\in(0,1)$, and there exists $\gamma\in\mathbb{Z}$, $\mathcal{F}\subset S^{S\times U\times W}$, and
@@ 3.1 The main algorithm, theorem, and proof
Let $\phi_{f}:X\rightarrow\mathbb{B}$ be the failure check function (which composes the testing system dynamics (1), state trajectory collection, and the intersection check against $\mathcal{C}$) with a boolean output (w.l.o.g., let $\phi_{f}(x)\in\{0,1\},\forall x$). We have $r^{*}=\mathbb{E}_{p}\phi(x)$ and the risk estimate through Algorithm 2 as $r=\frac{1}{n}\sum_{i=1}^{n}\phi_{f}(x_{i})p(x_{i})/q(x_{i})$. Let $L=D(p\mid\mid q)$. Then, for $n=e^{L+t}$ with some $t\in\mathbb{R}_{\geq 0}$,
@@
$\leq\sqrt{\mathbb{E}(\phi_{f}(x)^{2})}\Bigg{(}e^{-\frac{c}{4}}+2\sqrt{\mathbb{P}\Big{(}\log\frac{p(x)}{q(x)}>L+\frac{c}{2}\Big{)}}\Bigg{)}$
$\leq\bar{r}\Bigg{(}e^{-\frac{c}{4}}+2\sqrt{\mathbb{P}\Big{(}\log\frac{p(x)}{q(x)}>L+\frac{c}{2}\Big{)}}\Bigg{)}\leq\tau^{\prime}.$
@@
With $|r-r^{*}|\leq 1$, and by Hoeffding’s inequality, we further have $\mathbb{P}(|r-r^{*}|\geq\tau)\leq 2e^{-2n\tau^{2}}$. This proves the $\beta$-repeatability of estimation accuracy $\tau$ except with probability $2e^{-2n\tau^{2}}$. Furthermore, for $\gamma$-reliability to be satisfied, the number of samples $n$ needs to be upper bounded by $\gamma$. It is thus expected that the importance distribution $q$ satisfying
@@ 3.2 Discussions
Moreover, the nominal and importance distributions are $p$ and $q$, respectively. Note in the general statistics applications, $p$ and $q$ may not be known properties [4], hence the presented results may not be as useful or applicable. However, in the context of testing, $p$ is often expected to be provided as part of the specifications made by standard organizations and regulatory authorities, while $q$ is required for testing execution by operators and testing agencies.
@@
Not every importance sampling based testing algorithm is an accelerated testing algorithm. The accelerated performance largely relies on the properly chosen importance distribution $q$, which is out of the scope of this paper [35, 34, 12, 15, 14]. The results in this section are applicable for an arbitrary pair of target and importance distributions $(p,q)$.
@@
The reliability of Algorithm 3, revealed by Theorem 3.1, is jointly determined by the Kullback–Leibler (KL) divergence between the two distributions and the value of $c$ which is mainly associated with the typical order of fluctuations of $\log(p(X)/q(X))$ from its expected value. In practice, the KL-divergence between $p$ and $q$ is immediate given the above, and one can numerically approximate the minimum estimate of $c$ that satisfies (4) (e.g., an incremental trial-and-error search of $c$ values from zero as demonstrated in line 5-6 in Algorithm 3 or some other bisection alternatives). Note $c_{r}$ is preferably a small value to enhance the efficiency, but the search with large steps will not undermine the theoretical properties dictated from Theorem 3.1. These two values (order of fluctuations and KL-divergence) critically influence the necessary testing efforts to minimize the risk estimation error (subject to the $\ell_{1}$-norm bound of $\tau$) with a high probability of $1-2e^{-2n\tau^{2}}$, thereby insuring $\gamma$-reliability (and $\beta$-repeatability).